OpenStudy (anonymous):

I just need to make sure I did this properly. I am supposed to write this as a single logarithm: 2/3(4 ln x-5 ln y) and I re-wrote it as ln(4x/5y)^2/3 is that right?

OpenStudy (anonymous):

I am going to re-write these using the equation editor just so the question is more clear... one second.

OpenStudy (anonymous):

\[2/3(4 \ln x - 5\ln y) does that equal \ln (4x/5y)^{2/3}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

should be ln(x^4/y^5) inside

OpenStudy (anonymous):

coefficients become exponents

OpenStudy (anonymous):

ok, what about the 2/3?

OpenStudy (anonymous):

that part you did correct

OpenStudy (anonymous):

lol.... ok, thanks for the help!